OpenStudy (anonymous):

26 students, each girl drew a triangle and each boy drew a rectangle. If there were 92 sides drawen by the students, how many girls were in the class?

OpenStudy (radar):

Let x = number of girls 26-x = number of boys 3x = number of lines the girls drew 4(26-x)= number of lines the boys drew. Now the sum is equal to 92 Do you need any more guidance, or can you take it from there?

OpenStudy (radar):

Don't look at this until you have at least tried to solve it. 3x+4(26-x)=92 3x+104-4x=92 3x-4x=92-104 -x=-12 x=12 (12 girls) 26-12= 14 (14 boys)